# GeneMamba: Efficient and Effective Foundation Model on Single Cell Data

### v1

@@ 5 Conclusion
In this study, we present GeneMamba, a foundational model designed to advance single-cell analysis. By treating cells as sentences and genes as tokens, and with the BiMamba block and biological aware loss function, GeneMamba achieves a deeper understanding of cellular heterogeneity and gene relationships. This work underscores the potential of adapting large language models for biological data analysis, bridging computational innovation with biological insights. Future directions include enhancing the model with mechanisms like cross-attention for improved transfer learning, scaling to even larger datasets, and creating accessible tools for the research community.
Code Availability
Datasets and code are available at the github address: https://github.com/MineSelf2016/GeneMamba
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our work has potential societal implications, but none warrant specific emphasis here.
@@
Appendix A Dataset Construction
@@ B.1 Downstream Datasets
The original Myeloid dataset is highly imbalanced. To better evaluate the models’ performance, we excluded rare cell types (proportion $<$ 0.05) to create the Myeloid_b dataset. Figure 6 illustrates the cell type distribution in both the original and modified Myeloid datasets.
@@ B.2 Downstream Tasks
We standardize the downstream datasets by converting the batch key to ”batch” and the label key to ”celltype.” We then use a manual stratified split to divide the dataset into train/test with a 0.9/0.1 ratio to avoid missing rare classes in the test set. For example, in the COVID dataset, using ‘train_test_split()‘ directly would cause the 22nd class to be missing, as it contains only 3 samples, and a 0.3 test sample rate would result in zero samples, causing errors during AUC-ROC calculation. The split column is labeled ”partition,” where ”train” is for training and ”test” is for testing.
@@
Figure 7: Gene Rank Reconstruct framework of GeneMamba Gene Imputation framework evaluating rank reconstruction by predicting (n+1)-th tokens iteratively.
@@
In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper (Du et al., 2019). And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair. We evaluate how the distribution of positive pairs and negative pairs differ by calculating the eucilean distance, KL-Divergence, and JS-Divergence of the two distributions, the result is shown in figure 5.
@@ B.3 Evaluation Metrics
  • GeneFormer (Theodoris et al., 2023): A transformer-based model specifically designed for single-cell analysis. It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
  • scGPT (Cui et al., 2024a): A generative pre-trained transformer adapted for single-cell data analysis. scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
  • scFoundation (Hao et al., 2024): A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles. It excels in tasks requiring integration across batches and datasets.
  • scBERT (Yang et al., 2022): A BERT-like architecture adapted for single-cell data. scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
  • Harmony (Korsunsky et al., 2019): A batch-effect correction method that aligns single-cell datasets from different conditions or experiments. This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@
Appendix C More Related Work
C.1 Discretization of Gene Expression
@@
Fixed-Size Binning: As seen in scBERT, fixed-size binning divides the range of expression values into equally sized intervals. Genes falling within the same interval are represented identically. This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin (Boiarsky et al., 2023)), masking subtle but biologically relevant differences.
@@
Value Projection Strategy: The value projection strategy, first introduced by TOSICA (Chen et al., 2023), has since been adopted by scCLIP (Xiong et al., 2023), SpaFormer (Wen et al., 2023b), CellLM (Zhao et al., 2023), and CellPLM (Wen et al., 2023a). In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations. This method retains the full resolution of the data, avoiding the information loss associated with binning. However, feeding continuous values into transformer-based models diverges from traditional tokenization approaches in NLP, potentially impacting model performance. Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
@@
Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK (Shen et al., 2022) for single-cell transformers. This approach ranks gene expression values within each cell and retains only the top-ranked genes. The remaining genes are either truncated or represented as a lower-priority group. By focusing on the top-ranked genes, this strategy reduces sequence length, significantly improving computational efficiency. It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
C.2 Model Architectures for Single-Cell Analysis
@@
Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA (Yu et al., 2024), a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena (Oh et al., 2023), a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene (Gong et al., 2024b), an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort (Jiao et al., 2023), a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer (Fan et al., 2024), which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT (Nouri, 2025), a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer(Cui et al., 2024b), a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS (Xu et al., 2023), an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT (Dong et al., 2024), a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST (Zhao et al., 2024a), an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis. These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
@@
Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures (Boiarsky et al., 2023; White & Harris, 2024). While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy. These limitations become particularly evident when processing the ultra-long sequences typical of single-cell RNA-seq data. To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency. For example, scBERT integrates Performer (Choromanski et al., 2020) to approximate attention calculations, while scGPT employs Flash Attention (Dao et al., 2022) to accelerate training by reducing memory overhead. Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes. To address the intrinsic limitations of transformers, Bidirectional Mamba (Bi-Mamba) has been proposed as a novel alternative architecture. Bi-Mamba leverages state-space models (SSMs) to process ultra-long sequences with linear computational complexity in both time and memory. This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.
Appendix D Discussion of Scalability
@@
Mamba Layers Hidden Size Seq Length Number of Params Training Time (hours/m)
@@
Table 6 compares the computational efficiency of the Transformer and GeneMamba backbones. GeneMamba requires significantly fewer FLOPs per sample and achieves faster training times across all sequence lengths. For example, at a sequence length of 2048, GeneMamba’s training time is 0.1284 seconds/sample compared to the Transformer’s 0.2227 seconds/sample, a reduction of approximately 42%. This efficiency becomes even more pronounced at longer sequence lengths.
Model Backbone Parameters Forward + Backward FLOPs/sample Seq Length Training Time (seconds/sample)
@@
Table 6: Computational efficiency comparison of Transformer and GeneMamba backbones at different input sequence lengths. GeneMamba significantly reduces computational cost while maintaining high efficiency.
@@
Total FLOPs $\text{Total FLOPs required}=1.58\times 10^{12}\times 2\times 10^{6}=3.16\times 10^{18}\text{ FLOPs.}$.
@@
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{19.5\times 10^{12}\text{ FLOPs/sec}}\approx 1.621\times 10^{5}\text{ seconds}=45\text{ hours.}$.
@@
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{312\times 10^{12}\text{ FLOPs/sec}}\approx 1.013\times 10^{4}\text{ seconds}=2.8\text{ hours.}$.
@@
Appendix E Extended Figures
@@
Appendix F Extended Tables

### v2

@@ 5 Conclusion
We present GeneMamba, a scalable model for single-cell analysis that treats cells as sentences and genes as tokens. Using the BiMamba block and biologically informed losses, GeneMamba captures gene relationships and cell diversity effectively. It performs well across key tasks and offers a strong foundation for future single-cell studies.
Broader Impacts and Limitations. GeneMamba can help researchers analyze large-scale single-cell data, offering insights into gene function and cell behavior. This has potential benefits in areas like disease research and drug development. However, the model may struggle with rare cell types or subtle signals from low-expression genes. It also requires significant computing resources for pretraining, which could limit broader use. Future work could focus on improving efficiency and expanding to more flexible architectures.
@@
Appendix A Dataset Construction and Setup
@@ B.1 Downstream Datasets
The original Myeloid dataset is highly imbalanced. To better evaluate the models’ performance, we excluded rare cell types (proportion << 0.05) to create the Myeloid_b dataset. Figure 5 illustrates the cell type distribution in both the original and modified Myeloid datasets.
@@ B.2 Downstream Tasks
We standardize the downstream datasets by converting the batch key to "batch" and the label key to "celltype." We then use a manual stratified split to divide the dataset into train/test with a 0.9/0.1 ratio to avoid missing rare classes in the test set. For example, in the COVID dataset, using ‘train_test_split()‘ directly would cause the 22nd class to be missing, as it contains only 3 samples, and a 0.3 test sample rate would result in zero samples, causing errors during AUC-ROC calculation. The split column is labeled "partition," where "train" is for training and "test" is for testing.
@@
Figure 6: Gene Rank Reconstruct framework of GeneMamba
@@
In this experiment, we evaluate the GeneMamba’s performance to distinguish between positive pairs and negative pairs, the reference data is collected in the Gene2Vec paper [du2019gene2vec]. And we generate the embeddings using different models for positive gene pairs and negative gene pairs, then we measure the embeddings similarity of each pair. We evaluate how the distribution of positive pairs and negative pairs differ by calculating the eucilean distance, KL-Divergence, and JS-Divergence of the two distributions, the result is shown in Figure 7.
@@ B.3 Evaluation Metrics
  • GeneFormer [theodoris2023transfer]: A transformer-based model specifically designed for single-cell analysis. It uses a gene-token representation to capture complex gene-gene relationships and downstream cellular heterogeneity.
  • scGPT [cui2024scgpt]: A generative pre-trained transformer adapted for single-cell data analysis. scGPT utilizes a pretraining-finetuning paradigm to handle various single-cell tasks, including cell-type classification and clustering.
  • scFoundation [hao2024large]: A foundational model for single-cell data analysis that incorporates self-supervised learning techniques to generate robust embeddings for single-cell profiles. It excels in tasks requiring integration across batches and datasets.
  • scBERT [yang2022scbert]: A BERT-like architecture adapted for single-cell data. scBERT leverages bidirectional contextual embeddings to model complex interactions between genes within cells, making it suitable for both classification and clustering tasks.
  • Harmony [korsunsky2019fast]: A batch-effect correction method that aligns single-cell datasets from different conditions or experiments. This method is often used as a preprocessing step to ensure integration and compatibility across datasets before applying downstream machine learning methods.
@@
Appendix C Gene Correlation Analysis
Differentiating Positive and Negative Gene Pairs In the first experiment, GeneMamba’s ability to differentiate between "positive" and "negative" gene pairs was assessed using Gene2Vec embeddings. Nearly 30,000 gene pairs, evenly distributed as positive (1) or negative (0), were analyzed for similarity scores using cosine similarity and Pearson correlation. GeneMamba demonstrated a significant separation between the mean similarity scores of positive and negative pairs (Figure 7 a, b). Compared to baseline models, GeneMamba achieved a clearer distinction, emphasizing its capacity to accurately represent gene relationships.
Embedding Alignment To quantify discrimination power, we compared GeneMamba’s embeddings with those of scGPT and scFoundation. The "distance" between positive and negative pair distributions was measured using Euclidean distance, KL divergence, and JS divergence. The result is shown in Appendix Figure 15. GeneMamba consistently outperformed the other models across all metrics, highlighting its superior ability to distinguish between gene pair types and reinforcing its robustness in modeling gene relationships.
Gene-Gene Topology Analysis The ability to capture gene-gene topologies was analyzed using adjacency matrices derived from embeddings. For the PBMC12K dataset, distances between embeddings of randomly selected gene pairs (100 cells in our case) were calculated and binarized based on threshold value. Then we build the topology graph based on the adjacent matrix (Figure  7c) for each evaluated model. Notably, while identifying unique model-specific topologies, all three models captured biologically significant genes and their interactions, such as RALA, DECR1, and CBL1. This suggests that the foundation models share a similar representation space.
Appendix D Perturbation Analysis
To more comprehensively evaluate the generalizability and biological relevance of GeneMamba, we conducted a perturbation prediction study using protocols established by scGPT [cui2024scgpt]. Perturbation studies are critical for assessing a model’s ability to simulate cellular responses to genetic interventions—a fundamental requirement in functional genomics and drug discovery. We employed the Perturb-seq dataset from Norman [norman2019exploring], which include two-gene combinatorial perturbations. Specifically, we compared Mamba-based GEARS embeddings with the original GEARS model (baseline), following the benchmarking criteria previously proposed.
Figure 8: Mean Squared Error of DE analysis
Since single-cell-level ground truth was unavailable, we evaluated performance using the averaged Mean Squared Error (MSE) of the top 20 differentially expressed (DE) genes before and after perturbation, and the result is shown in Figure 8. GeneMamba achieved lower MSE across these genes compared to the baseline, suggesting improved fidelity in modeling transcriptional changes. Additionally, we quantified prediction accuracy by measuring the proportion of top 20 DE genes whose predicted expression fell within the 45–55% quantile range of the true expression distribution, where the Mamba-based model again showed superior alignment with ground truth.
Figure 9: ETS2 + CEBPE perturbation result
Figure 10: Magnitude scores computed for all test perturbing combinations on the Norman dataset.
Furthermore, in a representative combinatorial perturbation case (ETS2 + CEBPE), as shown in Figure 9, GeneMamba-generated predictions (red) closely matched the ground truth (purple) and outperformed the baseline (blue) in capturing post-perturbation expression shifts. Correlation analysis between predicted and ground-truth magnitude scores across all test perturbations yielded a higher Pearson coefficient (r = 0.5686, p < 1e-6) for GeneMamba, compared to the baseline GEARS (r = 0.4117), as shown in Figure 10. Lastly, the identification of suppressor and synergistic gene interactions revealed broader and more accurate overlap with the verified perturbation set using GeneMamba. These findings underscore the model’s enhanced interpretability and robustness under gene perturbation scenarios.
Appendix E More Related Work
E.1 Discretization of Gene Expression
@@
Fixed-Size Binning: As seen in scBERT, fixed-size binning divides the range of expression values into equally sized intervals. Genes falling within the same interval are represented identically. This method, however, often leads to significant information loss, as most genes with low expression levels cluster within the same bin (e.g., the [0,1) bin [boiarsky2023deep]), masking subtle but biologically relevant differences.
@@
Value Projection Strategy: The value projection strategy, first introduced by TOSICA [chen2023transformer], has since been adopted by scCLIP [xiong2023scclip], SpaFormer [wen2023single], CellLM [zhao2023large], and CellPLM [wen2023cellplm]. In this approach, continuous gene expression values are directly projected into high-dimensional embeddings through linear or non-linear transformations. This method retains the full resolution of the data, avoiding the information loss associated with binning. However, feeding continuous values into transformer-based models diverges from traditional tokenization approaches in NLP, potentially impacting model performance. Additionally, continuous embeddings may be more sensitive to noise and batch effects, necessitating robust preprocessing steps.
@@
Rank-Based Strategy: The rank-based strategy was first introduced in iSEEEK [shen2022universal] for single-cell transformers. This approach ranks gene expression values within each cell and retains only the top-ranked genes. The remaining genes are either truncated or represented as a lower-priority group. By focusing on the top-ranked genes, this strategy reduces sequence length, significantly improving computational efficiency. It also aligns well with the biological prioritization of highly expressed genes in specific cell states or conditions.
E.2 Model Architectures for Single-Cell Analysis
@@
Recent advancements in model architectures have addressed the computational challenges of transformer-based models in single-cell RNA sequencing (scRNA-seq) analysis, with notable innovations including scTCA [scTCA], a hybrid Transformer-CNN architecture designed for imputation and denoising of single-cell read counts to enhance downstream analyses; scHyena [scHyena], a foundation model for full-length scRNA-seq analysis in brain tissues that uses a novel transformer architecture to process data without information loss, improving tasks like cell type classification and data imputation; xTrimoGene [gong2024xtrimogene], an efficient and scalable representation learner employing an asymmetric encoder-decoder transformer architecture to reduce computational requirements while maintaining high accuracy in tasks such as cell type annotation and drug combination prediction; scTransSort [scTransSort], a transformer-based model for intelligent cell type annotation that leverages self-attention mechanisms to extract features from scRNA-seq data, achieving high accuracy and performance; scGraphformer [scGraphformer], which employs transformer-based graph neural networks to provide accurate and scalable cell type annotations, uncovering cellular heterogeneity and interactions in scRNA-seq data; scGFT [scGFT], a train-free, cell-centric generative model adept at synthesizing scRNA-seq data to enhance data augmentation and analysis; White-Box Diffusion Transformer[WhiteBoxDiffusionTransformer], a hybrid model combining diffusion models and white-box transformers to generate synthetic and biologically plausible scRNA-seq data, improving training efficiency and resource utilization; STGRNS [STGRNS], an interpretable transformer-based method for inferring gene regulatory networks from scRNA-seq data to enhance understanding of gene interactions; scRDiT [scRDiT], a generative approach using diffusion transformers to create virtual scRNA-seq data, facilitating the study of gene expression dynamics; and TransformerST [TransformerST], an unsupervised model based on the transformer architecture for super-resolution in spatial transcriptomics, improving spatial gene expression analysis. These models collectively advance the field by addressing key challenges in scRNA-seq data processing and interpretation.
@@
Traditional single-cell large language models (LLMs) are predominantly based on transformer architectures [boiarsky2023deep, white2024comprehensive]. While transformers have demonstrated exceptional performance in various single-cell transcriptomics tasks, their high computational and memory requirements make training large models time-intensive and resource-heavy. These limitations become particularly evident when processing the ultra-long sequences typical of single-cell RNA-seq data. To alleviate these challenges, some transformer-based models have incorporated optimizations to improve efficiency. For example, scBERT integrates Performer [choromanski2020rethinking] to approximate attention calculations, while scGPT employs Flash Attention [dao2022flashattention] to accelerate training by reducing memory overhead. Despite these advancements, the overall time required to train such models remains substantial, especially for datasets with millions of cells and thousands of genes. To address the intrinsic limitations of transformers, Bidirectional Mamba (Bi-Mamba) has been proposed as a novel alternative architecture. Bi-Mamba leverages state-space models (SSMs) to process ultra-long sequences with linear computational complexity in both time and memory. This represents a significant departure from traditional transformer-based methods, offering a scalable and efficient solution for the high-dimensional, high-throughput nature of single-cell transcriptomics data.
Appendix F Discussion of Scalability
@@
Mamba Layers Hidden Size Seq Length Param Size Training Time (h/million)
@@
Table 6 compares the computational efficiency of the Transformer and GeneMamba backbones. GeneMamba requires significantly fewer FLOPs per sample and achieves faster training times across all sequence lengths. For example, at a sequence length of 2048, GeneMamba’s training time is 0.1284 seconds/sample compared to the Transformer’s 0.2227 seconds/sample, a reduction of approximately 42%. This efficiency becomes even more pronounced at longer sequence lengths. These results confirm that our conclusions remain valid: FlashAttention only optimizes memory access but keeps $O(L^{2})$ scaling, whereas GeneMamba achieves true linear-time scaling and much lower computational cost for long sequences.
Table 6: Comparison of computational efficiency across Transformer, FlashAttention, and GeneMamba.
Model Backbone Parameters FLOPs/sample Seq Length Training Time (s/sample)
@@
$\text{Total FLOPs required}=1.58\times 10^{12}\times 2\times 10^{6}=3.16\times 10^{18}\text{ FLOPs.}$
@@
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{19.5\times 10^{12}\text{ FLOPs/sec}}\approx 1.621\times 10^{5}\text{ seconds}=45\text{ hours.}$
@@
$t=\frac{3.16\times 10^{18}\text{ FLOPs}}{312\times 10^{12}\text{ FLOPs/sec}}\approx 1.013\times 10^{4}\text{ seconds}=2.8\text{ hours.}$
@@
Appendix G Extended Figures
@@
Appendix H Extended Tables
@@
We present the complete baseline results directly below, including both Accuracy and Macro-F1 scores, along with the detailed experimental configurations used to obtain them.
Table 8: Baseline results comparing HVG, Seurat, scVI, scANVI, and GeneMamba. Bold numbers denote best performance.
Dataset Metric HVG Seurat scVI scANVI GeneMamba
hPancreas Accuracy 0.9770 0.9668 0.9597 0.9662 0.9713
Macro-F1 0.7029 0.7346 0.6631 0.7023 0.7710
MS Accuracy 0.7055 0.8600 0.7790 0.8675 0.6825
Macro-F1 0.5403 0.7335 0.6357 0.7424 0.5342
Myeloid Accuracy 0.6052 0.6571 0.6029 0.6332 0.6607
Macro-F1 0.3334 0.3627 0.3330 0.3400 0.3650
Myeloid_b Accuracy 0.8942 0.9018 0.8929 0.9231 0.9603
Macro-F1 0.8825 0.8930 0.8780 0.9147 0.9235
Experimental Setup
  • HVG (Highly Variable Genes): We selected the top 2,000 highly variable genes per dataset to reduce dimensionality while retaining biologically informative variation.
  • Seurat (R v4.4.0):
    1. 1.
      Standard PCA for dimensionality reduction.
    2. 2.
      Random Forest as backbone module.
    3. 3.
      Clustering resolution was tuned to align with reference labels for fair comparison.
  • scVI / scANVI (scvi-tools==1.1.0):
    1. 1.
      Latent space dimension: 32
    2. 2.
      Classification head: 2-layer MLP
    3. 3.
      Activation: ReLU
    4. 4.
      scANVI initialized from scVI embeddings with default semi-supervised ratio.
  • Training:
    1. 1.
      100 epochs
    2. 2.
      Batch size: 512
    3. 3.
      Adam optimizer, learning rate: 1e-3
    4. 4.
      Early stopping was monitored on validation loss.
Analysis and Observations
  1. 1.
    Overall performance: GeneMamba achieves the best performance on 3 of the 4 datasets and shows particularly strong Macro-F1 scores, indicating superior recognition of minority or rare cell populations.
  2. 2.
    Macro-F1 vs Accuracy:
    • scANVI achieves competitive or slightly higher Accuracy on some datasets (e.g., MS) but falls short in Macro-F1, implying its predictions are skewed toward majority classes.
    • GeneMamba’s higher Macro-F1 demonstrates balanced prediction across diverse cell types.
  3. 3.
    Biological implications: Balanced performance is crucial in single-cell analysis, as rare populations often carry key biological insights. GeneMamba’s results reflect its strength in capturing both majority and minority cell behaviors.
Below, we present the results across four benchmark datasets, comparing GeneMamba with six strong baselines:
Cell Type Annotation Performance (Expanded)
Table 9: Expanded benchmarks comparing GeneMamba with six strong baselines. Bold values denote the best performance.
Dataset Metric GeneFormer scGPT scFoundation scBERT scMulan GeneMamba
hPancreas Accuracy 0.9665 0.9710 0.9602 0.9594 0.9501 0.9713
Macro-F1 0.7450 0.7632 0.7101 0.7381 0.7012 0.7710
MS Accuracy 0.7650 0.8471 0.7763 0.8054 0.7888 0.6825
Macro-F1 0.6220 0.6630 0.6812 0.6412 0.6450 0.5342
Myeloid Accuracy 0.6445 0.6341 0.6446 0.6527 0.6483 0.6607
Macro-F1 0.3600 0.3562 0.3646 0.3482 0.3590 0.3650
Myeloid_b Accuracy 0.9540 0.9421 0.9574 0.9451 0.9487 0.9603
Macro-F1 0.9380 0.9434 0.9569 0.9261 0.9340 0.9235
Interpretation: GeneMamba achieves the best performance in both Accuracy and Macro-F1 on hPancreas and Myeloid, two biologically diverse datasets. Even in more challenging or imbalanced datasets such as MS and Myeloid_b, it remains competitive. These additional results provide stronger empirical evidence that GeneMamba is a robust and effective foundation model, standing out among both classical transformer-based and modern state space models.
Note: Results for scBERT and scMulan were obtained by fine-tuning their publicly released pretrained models on our standardized downstream annotation datasets, using the same MLP head, learning rate, and stratified train/test split as described in Appendix B.2.